# Large Language Models for Cryptocurrency Transaction Analysis: A Bitcoin Case Study

### v2

@@ 4.2. Level 1 - Foundational Metrics
    Response metrics. To evaluate if the responses from LLMs are correctly structured, struct_correctness is applied.
  • $\bullet$
@@
    Global metrics. To assess the ability of LLMs to basic metrics understanding for the entire transaction graphs, LLMs need to find the node with the largest in/out-degree (global_in/out_
    degree
    ), the node with the largest in/out-value (global_in/out_
    value
    ), and the node with the largest difference between input and output values/degrees (global_diff_degree/value).
  • $\bullet$
@@
    Node metrics. To investigate the capability of LLMs to understand foundational information of concrete nodes in transaction graphs, LLMs need to obtain the node’s in/out-degree (node_in/out_degree), the node’s in/out-value (node_in/out_
    value
    ), and the node’s special information (node_special_info_
    a/t
    ). The special information for the address node is time interval; while for the transaction node is if a specific node exists in the input/output node sets.
@@
Our results. Table 1 demonstrates that LLMs are excellent at node metrics. Accuracy for most metrics is between 98.50% and 100.00%. node_special_info_t is the exception. This may be due to the limited capability of LLMs to match many structurally similar data in transaction graphs. Compared with node metrics, however, for the global metrics, the accuracy significantly drops, ranging from 24.00% to 58.00%. Especially, we find that compared with the other metrics (35.00% to 58.00%) in global metrics, difference-related metrics are relatively low (24.00% to 34.00%). The reason for this may be that the capability of LLMs to calculate or compare is limited.
For different LLMs, i.e., GPT-4 and GPT4o, the most various points are struct_correctness in response metrics and global_in
/out_value
in global metrics. The enhancement of struct_correct
ness
@@
represents that the update and optimization of the LLMs improve the quality of the response format, which completely follows the requirements in prompts. Likewise, slightly improved global_in/out_value also illustrate the effectiveness of model upgrade. However, most of the metrics remain at similar levels, which shows the inherent flaws of LLMs for basic information understanding, especially global metrics, in transaction graphs.
Level 1 findings. For foundational metrics of transaction graphs, LLMs show high utility (98.50% to 100.00%) in node metrics and acceptable accuracy in several global metrics, especially for GPT-4o (44.00% to 58.00%). This illustrates that LLMs are excellent at obtaining specific node information in Bitcoin transaction graphs, but average at calculating or comparing data in the transaction graph.
4.3. Level 2 - Characteristic Overview
To measure LLMs’ capability in characteristic overview, we inquire about the top two most distinctive features of the subgraph. To achieve this, we randomly choose 16 subgraphs (two for each type) as knowledge pre-input to LLMs. Then, we randomly select 40 subgraphs (five for each type), and for each chosen subgraph, we combine it with the 16 reference subgraphs before inputting them into LLMs. Labels may affect the LLMs’ output, so we removed them from subgraphs.
We divide the quality of LLMs’ responses into three levels: high, average, and low. High-quality responses exclude invalid, inaccurate, or incorrect information. Average-quality responses include invalid or inaccuracies but no incorrect information. Low-quality responses contain incorrect information. In this context, an invalid response is one that, although accurate, fails to provide useful information. Apart from invalid/incorrect outputs, other responses generated by LLMs will be considered meaningful for analysis.
@@ 4.3. Level 2 - Characteristic Overview
Our results. Based on our comprehensive empirical analysis of selected samples, the percentages of high-quality, average-quality, and low-quality cases for GPT-4 are 62.50%, 26.25% (inaccurate and invalid are 7.50% and 18.75%), and 11.25%; for GPT-4o are 82.50%, 13.75% (inaccurate and invalid are 12.50% and 1.25%), and 3.75%. That means the proportions of meaningful responses for GPT-4 and GPT-4o are 70.00% and 95.00%. We select two examples, i.e.,bc1qah111bc1qahe54yxl33clnwdtleuh4cw0fw4df62t0tnuk2 and bc1q4w222bc1q4w090gzj7m9hy39l8e48gzz4uvdfn2x9y9g70h, to ensure each category is included to further illustrate our results, combined with visualization by Gephi. We apply the ID of address node $n_{0}$ to represent the corresponding subgraph.
@@
Figure 2. Examples with features
For bc1qah, the responses provided by GPT-4 exhibit average or low quality (Appendix A). The first response correctly identifies that nodes $n_{1}$ and $n_{2}$ have high out-degrees (Fig. 2(a)), but it inaccurately describes their in-degrees. Although both nodes have higher in-degrees than many in the reference subgraph (typically in-degree is 1, while 4 and 3 for $n_{1}$ and $n_{2}$), their in-degrees are not particularly high within this graph. For example, node $n_{14}$ has an in-degree of 198, significantly exceeding that of $n_{1}$ and $n_{2}$. In the second response, most transactions have similar in-value and out-value; while the differences between in-value and out-value of $n_{1}$ and $n_{2}$ are both about 0.003, which is trival.
@@
For bc1q4w, the responses by GPT-4 are high-quality or average-quality. The first response is accurate and illustrates high out-degree of $n_{1}$ in Layer 1 (Fig.2(b)). Also, transaction nodes, including $n_{886}$ and $n_{1228}$ in Layer 3, have high out-degree that exceed 400. The second response is correct and demonstrates the difference between in-value and out-value. However, it is invalid since many transactions follow this pattern, which is ineffectual for transaction analysis.
@@
Level 2 findings. Despite some inaccuracies responses exist, LLMs extract numerous beneficial features for transaction analysis, achieving 70.00% and 95.00% for GPT-4 and GPT4o, respectively. This indicates that LLMs have an exceptional characteristic overview understanding capability, especially GPT-4o, highlighting their utility in characteristics identification within Bitcoin transaction graphs.
4.4. Level 3 - Contextual Interpretation
We implement two experiments to evaluate the models’ contextual interpretation capabilities. The first is based on graph features, while the second utilizes raw graphs. In both experiments, we apply the few-shot prompting strategy. This involves giving the LLMs with labeled subgraphs (or their features) as references and one unlabeled subgraph (or its features) for explainable classification in each iteration of LLM query.
Graph feature-based classification. For the graph feature-based classification tasks, we employ five labeled subgraphs per category, where each subgraph is represented by its starting Bitcoin address (i.e., $n_{0}$) and its corresponding features, along with 500 randomly chosen unlabeled subgraphs with features. Except for inputting the values of these features to LLMs, we also input the corresponding descriptions. We select the ten most important features in BABD (10375557, ) (Table 2). Our analysis focuses on the accuracy, macro precision, recall, and F1 scores of LLMs and their performance in each class.
The overall accuracy of LLMs (Fig.3(a)) is between 39.83% to 46.07%, which is not impressive, as well as the macro precision, macro recall, and macro F1 score. However, this is acceptable since the reference samples are significantly few. In contrast, the top-3 accuracy achieves high, ranging from 67.20% to 71.02%. This suggests that while achieving exact classification might be challenging due to the limited data, LLMs can achieve relatively high accuracy in identifying the correct classes within their top three predictions.
Beyond overall metrics, class-specific results are displayed in Fig.4. GPT-4 and GPT-4o show strong performance in mining pool, achieving 80.00% and 95.00% precisions; They also reach extremely high recalls for darknet market which are 98.46% and 96.92%. For the majority of classes, GPT-4 and GPT-4o show better performance, though in some classes GPT-3 has similar or higher performance (e.g., recall for money laundering and precision for darknet market).
Table 1. Capability of LLMs in Foundational Metrics
Metrics GPT-4 GPT-4o
struct_correctness 80.00% (40/50) 100.00% (50/50)
global_in_degree 50.00% (20/40) 44.00% (22/50)
global_out_degree 50.00% (20/40) 58.00% (29/50)
global_in_value 37.50% (15/40) 56.00% (28/50)
global_out_value 35.00% (14/40) 48.00% (24/50)
global_diff_degree 27.50% (11/40) 34.00% (17/50)
global_diff_value 27.50% (11/40) 24.00% (12/50)
node_in_degree 99.25% (397/400) 99.40% (497/500)
node_out_degree 100.00% (400/400) 99.80% (499/500)
node_in_value 98.50% (394/400) 99.00% (495/500)
node_out_value 98.50% (394/400) 99.00% (495/500)
node_special_info_a 99.08% (216/218) 100.00% (259/259)
node_special_info_t 70.88% (129/182) 67.63% (163/241)
Table 2. Summary of important Metrics in analysis
@@ 4.4. Level 3 - Contextual Interpretation
Overall, GPT-4o achieves the highest top-3 accuracy (71.02%) and recall (47.45%), while also outperforming GPT-3.5 and GPT-4 in macro precision. However, its accuracy remains lower than GPT-4, and its F1 score, while surpassing GPT-3.5, is slightly lower than GPT-4. Nevertheless, GPT-4o demonstrates better stability, with fewer categories having macro precision, recall, or F1 scores below 10%. These results suggest that despite updates leading to marginal improvements in graph feature-based classification, GPT-4o still offers mild but limited advantages over GPT-3.5 and GPT-4.
Raw graph-based classification. For the raw graph-based classification, we utilized three labeled subgraphs per category and 101 unlabeled subgraphs, due to higher computational costs.
GPT-3.5GPT-4GPT-4o0020202020404040406060606080808080
(a) Graph feature-based
GPT-4GPT-4o0020202020404040406060606080808080
(b) Raw graph-based
Figure 3. Classification via different piplines (x for LLM models, y for percentage (%); bars 1st-5th represents accuracy, top-3 accuracy, precision, recall and F1 score, respectively)
@@
The overall accuracy of GPT-4 (Fig.3(b)) is significantly lower than desired. In contrast, GPT-4o surpasses all feature-based classification tasks, achieving an accuracy of 50.49%. Additionally, the macro precision, recall, and F1 scores of GPT-4 on raw graph-based classification tasks fall below those of GPT-3.5, GPT-4, and GPT-4o on graph feature-based classification tasks. Conversely, the metrics for GPT-4o on graph-based classification tasks are notably higher than those on graph feature-based classification tasks, all exceeding 50.14%. As for top-3 accuracy, GPT-4 is still lower than average, while GPT-4o achieves similar performance on raw graph-based classification compared with those on feature-based.
For the specific class as shown in Figure 5, GPT-4o performs better than GPT-4 in almost all metrics, in some classes such as darknet market, the differences of F1 scores are even close to 80%. One remarkable exception is that both GPT-4 and GPT-4o recorded zero scores in all metrics on blackmail; However, for GPT-4, classes including gambling, and Ponzi scheme recorded zero scores across all metrics, reflecting a huge failure in identifying relevant instances.
To sum up, GPT-4o achieves significantly higher performance than GPT-4 across all metrics in raw graph-based tasks and outperforms it in most specific classes. This demonstrates that optimizing LLMs may effectively enhance raw graph-based classification tasks.
Extended with more models. As illustrated in Fig.6, in the context of a very small number of reference samples, LLMs based on graph features perform significantly better than support vector machine (SVM) and MLP. However, they are inferior compared to DT, RF, CatBoost, and GNN. For LLMs based on raw graphs, GPT-4 slightly outperforms SVM and MLP, with noticeable advantages in accuracy and precision, although it is at a low level.
Notably, GPT-4o, also using raw graphs, achieves much higher accuracy than other LLMs using graph features and is very close to DT, RF, CatBoost, and GNN. Specifically, its precision greatly exceeds that of these tree models. This is crucial for illegal address detection because misidentifying legitimate addresses as illegal can have severe consequences, such as account suspension and even fund freezing. Overall, the performance of GPT-4o analyzing sampled graphs is comparable to the performance of tree models processing features selected through crafty feature engineering.
Analysis of classification explanation. Unlike traditional results of classification tasks, LLM-based outcomes can include detailed explanations that may be valuable for further analysis. However, our initial investigation indicates that these explanations are not always accurate. We choose two near-accurate explanations for 1Bsjsa3331BsjsaHST2Qohs8ZHxNHeZ1UfWhtxoKHEN and 124mpe444124mpePGM2vEqcHUK96wQXur9vS7Vn7Kdj (Appendix A) from GPT-4o to show the potential of LLMs and how they utilize different types of data.
These examples illustrate that for feature-based classification, adequate quantity and quality samples with labels are required to improve classification consequences since the selected features are already processed. In contrast, for graph-based classification, the quality of samples themselves seems to be also essential. The unsuitable compression of the graphs leading to much information loss may negatively affect the results. Also, the chain of thought (CoT) prompting (wei2022chain, ) could be used for further explanations, with detailed reasoning and better classification interpretability.
Level 3 findings. LLMs exhibit strong capabilities in contextual interpretation, achieving high top-3 accuracy rates even with limited data sets. They provide plausible explanations, although these are not always completely accurate. Despite these strengths, the overall accuracy of LLMs in classification tasks is moderate, indicating significant room for improvement in their performance.
0020202020404040406060606080808080100100100100Precision
0020202020404040406060606080808080100100100100Recall
BlackmailMarketExchangeGamblingPoolTumblerLaunderingPonzi00202020204040404060606060F1 Score
@@
Figure 4. LLMs’ performance in contextual interpretation using graph features (x axis for category, y for rate (%); GPT-3.5 in blue bar, GPT-4 in red, GPT-4o in brown).
0020202020404040406060606080808080100100100100Precision
0020202020404040406060606080808080Recall
BlackmailMarketExchangeGamblingPoolTumblerLaunderingPonzi0020202020404040406060606080808080F1 Score
@@
Figure 5. LLMs’ performance in contextual interpretation using raw graphs (x axis for category, y for rate (%); GPT-3.5 in blue bar, GPT-4 in red).
5. Discussion
Addressing RQs. We summarise our answers to RQs.
  • $\bullet$
    Answer to RQ1. We proposed LLM4TG format. The format demonstrates significant superiority (Fig.7). When combined with CETraS, it transforms tasks that were previously almost impossible into achievable outcomes, which means the combination enables experiments at various levels of understanding.
  • $\bullet$
    Answer to RQ2. We proposed a three-level framework for better understanding transaction graphs. For each level, we have established clear standards and conducted experiments to validate their practical application. The results underscore the framework’s effectiveness in enhancing the analytical capabilities of LLMs in this domain.
  • $\bullet$
    Answer to RQ3. We experimentally demonstrate that the engineered graph feature-based analyses focus on pre-computed metrics, simplifying complexity. In contrast, raw graph data analysis directly interprets structural relationships, providing deeper but more challenging insights. GPT-4o performs better with raw graph data compared to other LLMs using engineered features, whereas GPT-4 performs the worst with raw graph data (Fig.6).
Token consumption for different graph formats. We studied the differences in token consumption among various graph representation formats under the byte pair encoding-based tokenizer cl100k_base, the default GPT-3.5 and GPT-4 tokenizer. We chose three well-defined formats, i.e., GEXF, GML, and GraphML, for their efficiency and flexibility in representing transaction graphs. These widely used formats offer diverse encoding styles while supporting complex attributes of nodes and edges.
As in Fig.7, token usages of these formats exhibit significantly steeper curves, i.e., they consume a substantially larger number of tokens with an increasing number of nodes. These graph formats have extreme syntactic noise (i.e., great redundancy exists when representing graph data), which would result in numerous unnecessary tokens for LLMs to consume. The maximum token limits for GPT-3.5 and GPT-4/4o (16,385 and 128,000) are indicated by horizontal lines, where token usages of these formats surpass the GPT-4/4o limit when the number of nodes of a single graph exceeds approximately 500 to 750. Therefore, these graph representation formats are insufficient for LLMs to analyze transaction graphs.
How LLMs benefit transaction graph analysis. We summarise three advantages based on our experience applying LLMs.
  • $\bullet$
    High accuracy with minimal data. Utilizing LLMs for cryptocurrency transaction analysis and monitoring only requires minimal data. LLMs can provide high top-3 accuracy results with explanations as well as excellent precision when using raw graph data. Many addresses play multiple roles, such as exchanges may be involved in money laundering. LLMs explanations could be efficient for experts to further analyze, especially for some illicit transaction types with only a few samples.
  • $\bullet$
    Contextual aids. The underlying human motivations behind different types of transactions vary significantly and can reveal behavioral patterns often overlooked by purely data-driven approaches. Traditional models typically focus on identifying data patterns, neglecting the subtle reasoning behind user actions. In contrast, LLMs offer more context-aware interpretations, leveraging their extensive training in human language and behavior to infer intent or reasoning that conventional models fail to capture. While the explanations provided by LLMs may not always be entirely accurate, they still offer insights by interpreting the intent behind transactions.
  • $\bullet$
    Graph understanding:. LLMs can extract node information and provide meaningful overviews of transaction graphs, identifying complex patterns and relationships effectively. Unlike traditional algorithms, which often struggle with complex, high-dimensional data and contextual information embedded within graphs, LLMs leverage their capacity for processing raw data and capturing intricate patterns, relationships, and anomalies.
GPT-3.5-fGPT-4-fGPT-4o-fGPT-4-gGPT-4o-gSVMMLPDTRFCatBoostGNN0.20.20.20.20.40.40.40.40.60.60.60.6AccuracyPrecisionRecallF1 Score

### v3

@@ V-B Level 1 - Foundational Metrics
  • Response Metrics. To evaluate if LLM responses are correctly structured, struct_correctness is applied.
@@
  • Global Metrics. To assess the ability of LLMs to basic metrics understanding for the entire transaction graphs, LLMs need to find the node with the largest in/out-degree (global_in/out_degree), the node with the largest in/out-value (global_in/out_value), and the node with the largest difference between input and output values/degrees (global_diff_degree/value).
@@
  • Node Metrics. To investigate the capability of LLMs to understand foundational information of concrete nodes in transaction graphs, LLMs need to obtain the node’s in/out-degree (node_in/out_degree), the node’s in/out-value (node_in/out_value), and the node’s special information (node_special_info_a/t). The special information for the address node is time interval; while for the transaction node is if a specific node exists in the input/output node sets.
TABLE I: LLM Capability on foundational metrics
Metrics GPT-4 GPT-4o
struct_correctness 80.00% 100.00%
global_in_degree 50.00% 44.00%
global_out_degree 50.00% 58.00%
global_in_value 37.50% 56.00%
global_out_value 35.00% 48.00%
global_diff_degree 27.50% 34.00%
global_diff_value 27.50% 24.00%
node_in_degree 99.25% 99.40%
node_out_degree 100.00% 99.80%
node_in_value 98.50% 99.00%
node_out_value 98.50% 99.00%
node_special_info_a 99.08% 100.00%
node_special_info_t 70.88% 67.63%
@@
Results Analysis. Table I demonstrates that LLMs are excellent at node metrics. Accuracy for most metrics is between 98.50% and 100.00%. node_special_info_t is the exception. This may be due to the limited capability of LLMs to match many structurally similar data in transaction graphs. Compared with node metrics, however, for the global metrics, the accuracy significantly drops, ranging from 24.00% to 58.00%. Especially, we find that compared with the other metrics (35.00% to 58.00%) in global metrics, difference-related metrics are relatively low (24.00% to 34.00%). The reason for this may be that the capability of LLMs to calculate or compare is limited.
@@
For different LLMs, i.e., GPT-4 and GPT4o, the most various points are struct_correctness in response metrics and global_in/out_value in global metrics. The enhancement of struct_correctness represents that the update and optimization of the LLMs improve the quality of the response format, which completely follows the requirements in prompts. Likewise, slightly improved global_in/out_value also illustrate the effectiveness of model upgrade. However, most of the metrics remain at similar levels, which shows the inherent flaws of LLMs for basic information understanding in transaction graphs, especially global metrics.
Level 1 Findings. LLMs show consistently strong performance in node metrics, indicating a solid ability to extract explicit local information from transaction graphs. In contrast, performance on global metrics is noticeably weaker, especially for those requiring difference calculations. This disparity suggests that while models excel at retrieving localized details, they currently struggle with holistic graph reasoning and cross-node numerical comparisons.
V-C Level 2 - Characteristic Overview
To evaluate LLMs’ ability to identify distinctive subgraph characteristics, we provide representative reference subgraphs and test samples without labels to avoid bias. We divide the quality of LLMs’ responses into three quality levels:
  • High-quality responses exclude irrelevant, flawed, or incorrect information.
  • Average-quality responses include irrelevant or flawed but no incorrect information.
  • Low-quality responses contain incorrect information.
In this context, irrelevant responses are correct but do not answer the question or provide useful information; Flawed responses have small mistakes or misleading parts but are mostly right; Incorrect responses contain false or completely wrong information.
TABLE II: LLM capability on characteristic overview
Metric GPT-4 GPT-4o
High-quality 62.50% 82.50%
[…]
Flawed 7.50% 12.50%
Irrelevant 18.75% 1.25%
Low-quality 11.25% 3.75%
Meaningful 70.00% 95.00%
Overall Results. Table II shows that GPT-4o delivers substantially higher response quality than GPT-4. High-quality responses make up 82.50% of GPT-4o’s responses, compared with 62.50% for GPT-4. For average-quality responses, GPT-4o scores 13.75%, with 12.50% flawed and 1.25% irrelevant; While GPT-4 records 26.25%, with 7.50% flawed and 18.75% irrelevant. Low-quality responses are 3.75% for GPT-4o and 11.25% for GPT-4. Overall, GPT-4o produces meaningful outputs in 95.00% of cases versus 70.00% for GPT-4.
@@ V-C Level 2 - Characteristic Overview
Illustrative Examples. We select two examples bc1qah222bc1qahe54yxl33clnwdtleuh4cw0fw4df62t0tnuk2 and bc1q4w333bc1q4w090gzj7m9hy39l8e48gzz4uvdfn2x9y9g70h, to ensure each category is included to further illustrate our results, combined with visualization by Gephi. We apply the address identifier of node $n_{0}$ to represent the corresponding subgraph.
@@
Figure 2: Examples with features
Level 2 Example - bc1qah. GPT-4: $\bullet$ High in-degree and out-degree of transactions; Transactions like $n_{1}$ and $n_{2}$ have high in-degrees and out-degrees. [Average - Flawed] $\bullet$ Significant value differences; The in- and out-value of certain transactions show significant volume and discrepancies, such as in $n_{1}$ and $n_{2}$. [Low] GPT-4o: $\bullet$ High in-degree and out-degree of addresses in Layer 2; The addresses in Layer 2 have exceptionally high in-degree and out-degree values, indicating a large number of transactions. For example, address $n_{14}$ has an in-degree of 198 and out-degree of 188. [High] $\bullet$ High in-value and out-value in transactions; The transactions in this graph involve significant amounts of value. For instance, transaction $n_{1}$ has an in-value of 77.29740945 and out-value of 77.29452845. [High]
Level 2 Example - bc1q4w. GPT-4: $\bullet$ High out-degree in transaction nodes; A noticeable feature in this graph is the significant out-degree in some transaction nodes. For instance, $n_{1}$ has an out-degree of 600. [High] $\bullet$ Value discrepancy between in-value and out-value in transactions; The transaction node $n_{1}$ shows a discrepancy between the total incoming value (in-value: 27.69691553) and the outgoing value (out-value: 27.69184153). [Average - Irrelevant] GPT-4o: $\bullet$ High out-degree transactions; The transaction node $n_{1}$ has an extremely high out-degree of 600, meaning it distributes funds to 600 different address nodes. [High] $\bullet$ Large volume of transactions; The transaction node $n_{1}$ also handles a substantial volume of in-value and out-value (27.69691553). [High]
(i) Example - bc1qah. The responses provided by GPT-4 exhibit average or low quality. The first response correctly identifies that nodes $n_{1}$ and $n_{2}$ have high out-degrees (Fig. 2(a)), but it flawedly describes their in-degrees. Although both nodes have higher in-degrees than many in the reference subgraph (typically in-degree is 1, while 4 and 3 for $n_{1}$ and $n_{2}$), their in-degrees are not particularly high within this graph. For example, node $n_{14}$ has an in-degree of 198, significantly exceeding that of $n_{1}$ and $n_{2}$. In the second response, most transactions have similar in-value and out-value; while the differences between in-value and out-value of $n_{1}$ and $n_{2}$ are both about 0.003, which is trivial.
@@
(ii) Example - bc1q4w. The responses by GPT-4 are high-quality or average-quality. The first response is accurate and illustrates high out-degree of $n_{1}$ in Layer 1 (Fig.2(b)). Also, transaction nodes, including $n_{886}$ and $n_{1228}$ in Layer 3, have high out-degree that exceed 400. The second response is correct and demonstrates the difference between in-value and out-value. However, it is irrelevant since many transactions follow this pattern, which is ineffectual for transaction analysis.
@@
Level 2 Findings. Despite occasional irrelevant or incorrect responses, LLMs can extract informative structural and value-based features from Bitcoin transaction subgraphs without labels. GPT-4o shows clearer judgment in distinguishing genuinely significant features from trivial ones, and provides interpretations that reflect their analytical relevance; Whereas GPT-4 more often misjudges feature significance or emphasizes patterns of limited utility.
V-D Level 3 - Contextual Interpretation
We evaluate the LLMs’ contextual interpretation capability under two settings: one using graph features and the other using raw graphs. In both settings, we adopt a few-shot prompting strategy, where the model is given several labeled subgraphs (or their features) as references and one unlabeled subgraph for explainable classification in each iteration of LLM query.
GPT-3.5GPT-4GPT-4oDeepSeek02020404060608080AccuracyTop-3 Acc.PrecisionRecallF1
(a) Feature-based
GPT-4GPT-4oDeepSeek02020404060608080
(b) Raw graph-based
Figure 3: Classification via different LLMs (x for LLM models, y for percentage (%); bars 1st-5th represents accuracy, top-3 accuracy, precision, recall and F1 score, respectively)
Graph Feature-based Classification. In this setting, we provide LLMs with subgraph features and their textual descriptions, including both labeled and unlabeled samples. Specifically, we use the ten most important features identified in BABD [13] (Table III). We then evaluate the models’ performance across classes using accuracy, macro precision, recall, and F1 scores.
The overall accuracy of LLMs (Fig.3(a)) ranges from 39.83% to 46.68%, with similarly limited performance in macro precision, recall, and F1 scores. Although these results appear modest, they are partly attributable to the scarcity of reference samples, which constrains precise classification. Notably, the top-3 accuracy remains high, i.e., 67.20% to 72.43%, indicating that while exact class prediction is challenging under limited data, LLMs can still reliably narrow down the correct category within the top three candidates.
TABLE III: Summary of important metrics in analysis
@@ V-D Level 3 - Contextual Interpretation
Beyond the overall metrics, Fig. 4 reveals uneven, category-dependent strengths. GPT-4 and GPT-4o generally set the pace, pairing standout precision in mining pools at 80.00% and 95.00% with near-saturated recall on the darknet market at 98.46% and 96.92%. DeepSeek remains competitive where a recall-weighted balance matters, yielding the strongest F1 on pools and a slight edge on exchange. GPT-3.5, though weaker overall, leads on money laundering because all models share the same recall there and its precision is highest, producing the top F1. This suggests that LLM performance is complementary rather than absolute, with relative strengths distributed unevenly across categories in this context.
When comparing models, GPT-4o attains the highest recall of 47.45% and macro precision of 39.84%, suggesting greater consistency across categories. DeepSeek records the strongest overall accuracy at 46.68%, top-3 accuracy at 72.43%, and F1 score of 36.56%, though accompanied by notably lower precision of 33.22%. GPT-4 yields comparatively balanced precision and recall, whereas GPT-3.5 trails in most metrics. Taken together, these results show that performance advantages are distributed unevenly, with different models favoring different trade-offs.
@@
Focusing specifically on the GPT series, GPT-4o demonstrates better stability, with fewer categories having precision, recall, or F1 scores significantly low. These results suggest that despite updates leading to marginal improvements in graph feature-based classification, GPT-4o still offers mild but limited advantages over GPT-3.5 and GPT-4.
Raw Graph-based Classification. For raw graph-based classification, we rely on fewer subgraphs, as the higher computational cost limits the scale compared to feature-based setting.
@@
The overall accuracy of GPT-4 (Fig.3(b)) is significantly lower than desired. In contrast, GPT-4o surpasses all feature-based classification tasks, achieving an accuracy of 50.49%. The macro precision, recall, and F1 scores of GPT-4 on raw graph-based classification tasks fall below those of GPT-3.5, GPT-4, and GPT-4o on graph feature-based classification tasks. Conversely, the metrics for GPT-4o on graph-based classification tasks are notably higher than those on graph feature-based classification tasks, all exceeding 50.14%. As for top-3 accuracy, GPT-4 is still lower than average, while GPT-4o achieves similar performance on raw graph-based classification compared with those on feature-based. While DeepSeek achieves competitive accuracy of 41.58% and F1 score of 35.54% using raw graphs, clearly outperforming GPT-4 yet still trailing GPT-4o.
02020404060608080100100PrecisionGPT-3.5GPT-4GPT-4oDeepSeek
02020404060608080100100Recall
BlackmailMarketExchangeGamblingPoolTumblerLaunderingPonzi0202040406060F1 Score
@@
Figure 4: LLMs’ performance in contextual interpretation using graph features (x axis for category, y for rate (%); GPT-3.5 in blue bar, GPT-4 in red, GPT-4o in brown, DeepSeek in gray).
02020404060608080100100PrecisionGPT-4GPT-4oDeepSeek
02020404060608080100100Recall
BlackmailMarketExchangeGamblingPoolTumblerLaunderingPonzi02020404060608080100100F1 Score
@@
Figure 5: LLMs’ performance in contextual interpretation using raw graphs (x axis for category, y for rate (%); GPT-4 in red, GPT-4o in brown, DeepSeek in gray).
For the specific class as shown in Figure 5, GPT-4o performs better than GPT-4 in almost all metrics, in some classes such as darknet market, the differences of F1 scores are huge. One remarkable exception is that both GPT-4 and GPT-4o recorded zero scores in all metrics on blackmail; however, for GPT-4, classes including gambling and Ponzi scheme recorded zero scores across all metrics, reflecting a huge failure in identifying relevant instances. In contrast, DeepSeek achieves competitive performance in several categories, particularly darknet market, exchange, and gambling, where its F1 scores are markedly higher than GPT-4, although it still struggles in classes such as pool and Ponzi scheme, where GPT-4o maintains clear advantages.
To sum up, GPT-4o consistently achieves substantially higher performance than GPT-4 across all metrics in raw graph-based tasks and outperforms it in most specific classes. DeepSeek also delivers competitive results, showing notable strengths in categories such as darknet market and gambling, though it remains less stable than GPT-4o overall. These findings suggest that model optimization may enhance raw graph-based classification, while different LLM architectures may offer complementary advantages across categories.
GPT-3.5-fGPT-4-fGPT-4o-fGPT-4-gGPT-4o-gDeepSeek-fDeepSeek-gSVMMLPDTRFCatBoostGNN0204060AccuracyPrecisionRecallF1 Score